# RGD: Multi-LLM Based Agent Debugger via Refinement and Generation Guidance

### v1

@@ Abstract
Large Language Models (LLMs) have shown incredible potential in code generation tasks, and recent researches in prompt engineering have enhanced LLMs’ understanding of textual information. However, ensuring the accuracy of generated code often requires extensive testing and validation by programmers. While LLMs can typically generate code based on task descriptions, their accuracy remains limited, especially for complex tasks that require a deeper understanding of both the problem statement and the code generation process. This limitation is primarily due to the LLMs’ need to simultaneously comprehend text and generate syntactically and semantically correct code, without having the capability to automatically refine the code. In real-world software development, programmers rarely produce flawless code in a single attempt based on the task description alone, they rely on iterative feedback and debugging to refine their programs. Inspired by this process, we introduce a novel architecture of LLM-based agents for code generation and automatic debugging: Refinement and Guidance Debugging (RGD). The RGD framework is a multi-LLM-based agent debugger that leverages three distinct LLM agents—Guide Agent, Debug Agent, and Feedback Agent. RGD decomposes the code generation task into multiple steps, ensuring a clearer workflow and enabling iterative code refinement based on self-reflection and feedback. Experimental results demonstrate that RGD exhibits remarkable code generation capabilities, achieving state-of-the-art performance with a 9.8% improvement on the HumanEval dataset and a 16.2% improvement on the MBPP dataset compared to the state-of-the-art approaches and traditional direct prompting approaches. We highlight the effectiveness of the RGD framework in enhancing LLMs’ ability to generate and refine code autonomously.
@@ II Related Work
Large Language Models for Code Generation Recently, large language models (LLMs) have demonstrated outstanding performance multiple tasks especially in the code generation, showing great potential across various benchmarks [4] [15] [16]. However, when faced with complex problems, these models are prone to generating hallucinated responses [17]. Prompt engineering [18] has reduced the need for extensive fine-tuning of LLMs on specific downstream tasks, which can better understand user requirements through contextual information. Chain-of-thought prompting [6], which guides LLMs to provide step-by-step responses, enhances their reasoning abilities, leading to significant improvements over baselines on multiple benchmarks [19]. Tree-of-thought (ToT) [20] framework enhances the reasoning capabilities of large language models by structuring the reasoning process as a tree search. Each node represents a possible reasoning state, and edges denote transitions between states. In addition, researchers have explored enabling LLMs to debug the code they generate on their own [21], allowing the models to autonomously fix errors by understanding the causes from generated outputs [22] [8]. Beyond self-repair based on execution results, these frameworks can also leverage external tools for dynamic and automatic code correction [23].

### v2

@@ Abstract
Large Language Models (LLMs) have shown incredible potential in code generation tasks, and recent research in prompt engineering have enhanced LLMs’ understanding of textual information. However, ensuring the accuracy of generated code often requires extensive testing and validation by programmers. While LLMs can typically generate code based on task descriptions, their accuracy remains limited, especially for complex tasks that require a deeper understanding of both the problem statement and the code generation process. This limitation is primarily due to the LLMs’ need to simultaneously comprehend text and generate syntactically and semantically correct code, without having the capability to automatically refine the code. In real-world software development, programmers rarely produce flawless code in a single attempt based on the task description alone, they rely on iterative feedback and debugging to refine their programs. Inspired by this process, we introduce a novel architecture of LLM-based agents for code generation and automatic debugging: Refinement and Guidance Debugging (RGD). The RGD framework is a multi-LLM-based agent debugger that leverages three distinct LLM agents—Guide Agent, Debug Agent, and Feedback Agent. RGD decomposes the code generation task into multiple steps, ensuring a clearer workflow and enabling iterative code refinement based on self-reflection and feedback. Experimental results demonstrate that RGD exhibits remarkable code generation capabilities, achieving state-of-the-art performance with a 9.8% improvement on the HumanEval dataset and a 16.2% improvement on the MBPP dataset compared to the state-of-the-art approaches and traditional direct prompting approaches. We highlight the effectiveness of the RGD framework in enhancing LLMs’ ability to generate and refine code autonomously.
@@ II Related Work
Large Language Models for Code Generation Recently, large language models (LLMs) have demonstrated outstanding performance multiple tasks especially in the code generation, showing great potential across various benchmarks [4] [15] [16] [17]. However, when faced with complex problems, these models are prone to generating hallucinated responses [18]. Prompt engineering [19] has reduced the need for extensive fine-tuning of LLMs on specific downstream tasks, which can better understand user requirements through contextual information. Chain-of-thought prompting [6], which guides LLMs to provide step-by-step responses, enhances their reasoning abilities, leading to significant improvements over baselines on multiple benchmarks [20]. Tree-of-thought (ToT) [21] framework enhances the reasoning capabilities of large language models by structuring the reasoning process as a tree search. Each node represents a possible reasoning state, and edges denote transitions between states. In addition, researchers have explored enabling LLMs to debug the code they generate on their own [22], allowing the models to autonomously fix errors by understanding the causes from generated outputs [23] [8]. Beyond self-repair based on execution results, these frameworks can also leverage external tools for dynamic and automatic code correction [24] [25] [26].